OpenStudy (anonymous):

A rectangular barge 7 m long and 2 m wide floats in fresh water. How much will the barge sink when loaded with 600 kg of sand? Show all calculations leading to an answer.

OpenStudy (anonymous):

@sleepyjess @Michele_Laino

OpenStudy (michele_laino):

here we have to apply the Principle of Archimede

OpenStudy (anonymous):

I think the answer is 40mm"

OpenStudy (anonymous):

Water Density = 1,000kg/m^3 600/1,000= .6m^3 Area = 15m^2 V= (L)(W)(D)= .6^3 D= .6m^3/15m^2 = .04m = 40mm"

OpenStudy (michele_laino):

If I call with "h" the sinking level, then we can write: \[\Large LWh\delta g= Mg\] where: L is the length of the barge, W is the width of the barge \delta is the density of water g is gravity M is the mass of the loaded sand

OpenStudy (michele_laino):

is 600 Kg the mass of the sand ot its weight?

OpenStudy (michele_laino):

or*

OpenStudy (michele_laino):

I think that M=600 Kg is the mass of the sand

OpenStudy (anonymous):

yeah i think mass

OpenStudy (michele_laino):

so we get: \[\Large h = \frac{M}{{LW\delta }} = \frac{{600}}{{7 \times 2 \times 1000}} = ...\]

OpenStudy (michele_laino):

what is h?

OpenStudy (anonymous):

the depth of sinking?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

In mine i hade it D

OpenStudy (anonymous):

Made*

OpenStudy (michele_laino):

I got: \[\Large h \cong 43\;mm\]

OpenStudy (michele_laino):

so, you are right!

OpenStudy (anonymous):

Yay!

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

When a steadily-flowing gas flows from a larger-diameter pipe to a smaller-diameter pipe, what causes the pressure in the smaller pipe to drop? Use 3 – 4 complete sentences to explain your answer.

OpenStudy (anonymous):

do you know this? lol

OpenStudy (michele_laino):

if we can neglect all loss of energy of the gas, then we can apply the Principle of Bernoulli furtermore we can apply the equation of continuity

OpenStudy (anonymous):

@kaitlynmcurtis Bernoulli's principle states that the same V/s of fluid flowing will remain the same. If the radius of a pipe drops, then then the fluid has to travel faster through the pipe in order to maintain the same Volume per second... the only way the fluid can move faster is if there is a pressure drop from the larger pipe into the smaller pipe. Hope the helps!

OpenStudy (michele_laino):

for example, the equation of continuity is: \[\Large {A_1}{v_1} = {A_2}{v_2}\]  A is the cross sectional area v is the speed of the gas

OpenStudy (michele_laino):

so when the cross sectional area is minimum, then the speed of the gas is maximum

OpenStudy (michele_laino):

I have assumed that the density of our gas doesn't change as we go from section #1 to section #2

OpenStudy (michele_laino):

next the Principle of Bernoulli is: \[\Large gz + \frac{P}{\delta } + \frac{{{v^2}}}{2} = const\] where g is gravity, and z is the geodesic height

OpenStudy (anonymous):

I think I got this one done, could you guys help me with another? @cramos725 @Michele_Laino

OpenStudy (michele_laino):



OpenStudy (michele_laino):

ok!